# Quantum Computing in Wireless Communications and Networking: A Tutorial-cum-Survey

### v1

@@ III-D3 Applications of AI-driven QC for User Association
Several other studies [154][155] integrated AI techniques with QC algorithms within their solutions while considering network dynamics. For example, [154] discussed the path planning for cellular-connected UAVs to minimize the weighted sum of time cost and expected outage duration. Traditional offline optimization techniques face challenges due to practical considerations like local building distribution and directional antenna radiation patterns. To overcome this, the authors proposed a DRL solution that mapped the problem into a MDP, enabling the UAV to find the optimal flying direction within each time slot. However, the analysis did not establish direct correlations among multiple UAVs, simplifying the model into individual path planning for each UAV.
Due to robust capacity of QDRL for addressing challenges including those characterized by a large state and action spaces, it can also be effectively applied in the optimization of multiple objectives. [155] addressed multiple interconnected challenges, encompassing UAV trajectory planning, NOMA user grouping, and power allocation. Within this framework, users within a group maintained communication with the UAV within their communication ranges along the UAV trajectory. The decision regarding energy consumption by the UAV directly influenced the objective. Given the extensive and continuous nature of both action and state spaces, the paper introduced a QDRL approach to effectively tackle the complexity of the problem.
In conclusion, QC algorithms present a compelling approach for user association in wireless networks. Table VII offered a consolidated overview of these studies. However, there are two open issues in these QC algorithms addressing user association. First, these works typically relied on computer simulations for QC algorithm implementation. Second, few studies accounted for the impact of mobility on performance, even in UAV networks.
@@ IV-A Overview
This section explores how QC principles can revolutionize various aspects of networking, including routing, task offloading, and content caching as illustrated in Fig. 14. We aim to provide valuable insights into the current state of research, challenges, and future directions in leveraging QC to enhance networking capabilities.
@@ IV-B1 Overview
Routing presents fundamental challenges in network systems, encompassing the task of determining optimal paths for data packets to travel from a source to a destination. In wireless networks, where devices may be mobile and network conditions dynamically change, routing becomes even more complex. The primary goal of routing is to ensure efficient and reliable data transmission while minimizing latency, packet loss, and resource consumption. Existing approaches to routing encompass a variety of algorithms and protocols designed to address different aspects of routing optimization. Traditional methods like Optimized Link State Routing Protocol (OLSR) and Ad hoc On-Demand Distance Vector (AODV) attempt to dynamically establish and update routes based on network topology changes. However, they suffer from overhead due to frequent updates and incur latency in route discovery. Additionally, traditional routing algorithms may not adequately address challenges such as mobility, variable channel conditions, and energy constraints in wireless networks [157, 158].
QC-driven approaches offer promising solutions to overcome the limitations of traditional routing methods in wireless networks. By leveraging the principles of quantum mechanics, they can explore vast solution spaces more efficiently and potentially find optimal or near-optimal routes in significantly less time. In addition, QC enables the development of AI-driven routing algorithms that can adapt and learn from network dynamics, optimizing routing decisions in real-time. These quantum-enhanced approaches have the potential to improve routing efficiency, minimize latency, and enhance network reliability in wireless communication systems.
TABLE VIII: Summary of QC Works on Routing in Wireless Networks
@@
TABLE IX: Details of QC Algorithms in Routing
@@ IV-B2 Applications of Conventional QC for Routing
The paper [160] introduced a multi-phase method to address the vehicle routing with backhauls, mapping key elements of each phase onto QUBO models suitable for Ising machines. The two-phase method comprised a clustering phase followed by a routing phase, while the three-phase method included a depot clustering phase, a vehicle clustering phase, and a routing phase. The clustering-related phases were expansions of the knapsack problem, while the routing phases were tackled as TSP. Its process of modeling the problem as a QUBO formation can be regarded as a tutorial for applying QC in the field of networking.
@@
The paper [161] presented a multi-objective dynamic programming framework aimed at generating potentially Pareto-optimal routes by leveraging correlations among specific links within these routes, thereby significantly reducing the total number of routes. This framework extended the widely single-objective Bellman-Ford algorithm to handle multiple objectives. Additionally, the paper introduced a quantum-assisted algorithm that capitalized on both the dynamic programming framework and the synergies between Quantum Programming and Hybrid Programming to effectively solve the multi-objective routing problem. Similar works from the authors with different objectives or different QC algorithms were presented in [177] and [178]. The study [162], similar to [161], utilized the Boyer-Brassard-Hoyer-Tapp quantum search algorithm, but in a different scenario within self-organizing networks. They introduced an optimal multi-objective quantum-assisted algorithm aimed at assessing legitimate routes while incorporating the concept of Pareto optimality to reduce complexity. However, the study overlooked the influence of node mobility, a critical characteristic in self-organizing networks like Ad-hoc networks.
@@
The energy optimization model introduced in [163] targets sensor-enabled IoT environments, deriving energy consumption as the cost of energy-oriented paths. A QC-driven solution was proposed for optimizing the problem, emphasizing quantum representation, measurement, and rotation angle. However, a notable drawback of this approach was that minimizing the total energy cost across all relaying paths did not necessarily equate to prolonging the network lifetime. There were other works [164] based on quantum ant colony algorithms to solve routing problems. It utilized the metric, pheromone value inspired by ant colony algorithm, to measure a routing path. Due to dynamic network changes, it required a large computation time to track pheromone traits, especially in a large wireless sensor network.
@@
Additional popular quantum evolution based proposals for routing issues included QGA solutions [165, 166] and QPSO solutions [167, 168, 169]. [165] improved the network robustness of connectivity for IoT topology duo to node failures caused by malicious attacks. It proposed a QGA algorithm to maximize the cumulative sum of the number of nodes in the maximum connected sub-graph of the network. Additionally, [166] was designed to be resilient, which adapted to any changes in the planned topology of deployed sinks in anticipation of sensor failures. Also it achieved energy consumption balance among sensors by equalizing the path length and brings down the cost of wireless sensor networks by minimizing the number of sinks. Multicast routing in wireless mesh networks were considered with the objective of minimizing network delay [167]. The routing problem was converted into an integer programming problem with QoS constraints and was solved by QPSO combined with loop deletion operation. Connectivity in routing was explored in [168, 169]. [168] considered heterogeneous scenarios of network topology for optimal path configuration by exploring and exploiting the hunts. It applied multiple inputs from heterogeneous industrial IoT into QPSO techniques. The differential evolution operator and crossover operations were used for information interchange among the nodes to avoid trapping into local minima. Specifically, it studied the impact of p-degrees of network connectivity on network performance. The same scenario of IoT was also studied in [169] to optimize selection of cluster heads for multi-hop inter-cluster communication, which was different from user association in a single-cell communication in Subsection III-D. It proposed a cooperative multiple-input-multiple-output connectivity scheme between sensors and cluster heads to extend the network lifetime. QPSO was applied to select the optimum cooperative coalitions of each hop in the routing path.
@@
The preceding studies overlooked node mobility, a critical aspect in multi-hop wireless networks where devices and nodes dynamically moved in a network environment. Mobility plays an important role in various wireless communication scenarios, notably in 6G cellular networks, ad-hoc networks, vehicular networks, and UAV networks. Several works have employed QC algorithms to address the challenge. For instance, the optimized link state routing protocol, designed for mobile ad-hoc networks, was explored in [170] and [171] from the same researchers, both leveraging QGA approaches. The dynamic topology inherent in mobile ad-hoc networks pose significant challenges in finding and sustaining an optimal end-to-end path. However, the heuristic Q-learning strategy in RL enabled dynamic adjustment of routing paths through interactions with the surrounding environment. An enhanced QEA strategy has incorporated a novel Q-learning strategy, optimizing the selection of multi-point relays [170]. This improvement addressed the limitations of traditional protocols and demonstrates properties of convergence and global optimization. Other network aspects were additionally considered in [171], such as the delivery rate of data packets, and overhead of network topology control.
Additional networks, including UAV networks, were examined with a focus on node mobility in [172] and [173], utilizing QPSO and quantum ant lion algorithm, respectively. Specifically, [172] combined the differential evolution algorithm with QPSO to further enhance the performance of both algorithms. The primary objective was to describe a general route planning strategy. The route planning for UAV on the sea was formulated as an optimization problem. A route planner for UAV was designed to generate a safe and flyable path in the presence of different threat environments. [173] presented an approach for designing an energy-efficient neuro-fuzzy cluster-based topology construction with a metaheuristic route planning algorithm for UAVs. The model utilized three input parameters: residual energy in UAVs, average distance to nearby UAVs, and UAV degree, for selecting cluster heads and constructing the topology. Additionally, quantum ant lion based metaheuristic route planning was employed to select an optimal set of routes for inter-cluster UAV communication. However, the proposal divided the problem into two independent sub-problems, potentially leading to sub-optimal solutions.
@@ IV-B3 Applications of AI-driven QC for Routing
In recent years, AI techniques, such as RL and NNs, can be integrated with QC to enhance decision-making. AI-driven QC has emerged as a promising paradigm for addressing routing challenges to offer solutions in complex and dynamic wireless environments.
@@
A framework [174] was introduced aiming at leveraging QRL to enhance the design of next-generation wireless systems. The framework, termed QRL for relay and transmit power selection, enabled sensor nodes to make optimal decisions regarding relay and transmit power selection by leveraging both present and past local state knowledge. Initially, the relay and transmit power selection problem was formulated as MDP. Subsequently, the QRL optimization aspect of the MDP was addressed, focusing on jointly optimizing energy consumption and throughput to maximize network utility. To solve this MDP problem, a QRL algorithm was developed, drawing upon Grover’s iteration. The algorithm enabled efficient decision-making within the network, facilitating enhanced performance and resource utilization. Especially, the simulation was implemented using quantum virtual machine hosted on Rigetti’s Forest platform, which provided an additional approach to validating QC algorithms. Quantum Multiple Agent Reinforcement Learning (QMARL) was also applied in tackling routing problems. The work [175] introduced a quantum-based solution for MARL tailored for autonomous control and coordination of multiple robots within smart factory environments. The delivery problem shared similarities with the routing problem. It presented an innovative QMARL framework, featuring centralized training and decentralized execution. The framework leveraged parameter sharing on policy, design of variational quantum circuits, and 2-variables dense encoding. In particular, the proposed system enabled the QDRL server to make distributed and sequential decisions for each autonomous mobile robot. These decisions revolved around determining the optimal allocation of goods, including the number of goods to carry in each autonomous mobile robot and when to request quality control. The objective was to prevent both overflow and underflow scenarios during the delivery of goods by each autonomous mobile robot.
@@
The previous studies focusing on AI-driven QC overlooked the consideration of node mobility in wireless networks. This aspect was addressed by a limited number of studies, such as [176]. Multiple mobile UAV platforms acted as mobile BSs to gather data from other devices. Note that trajectory planning of multiple UAVs fell within the domain of routing, as data from devices were relayed by multiple UAVs to reach the destination. The proposed QMARL-based algorithm offered a notable advantage in reducing the action dimensionality. By employing project value measures, it effectively compressed the size into a logarithmic scale. This reduction proved beneficial for enhancing the convergence of training in large MARL scenarios.
@@ IV-C1 Overview
Research in edge computing task offloading encompasses a spectrum of crucial topics [179, 180]. Efficient task partitioning and scheduling, ensuring optimal allocation of tasks between edge devices and the cloud, form a foundational aspect. The effective management of limited resources on edge devices, spanning computing, storage, and communication, is a critical consideration [181].
TABLE X: Summary of QC Works on Task Offloading in Wireless Networks
@@
[184] A computing graph network with vertices as computing servers, and edges as vertex connection links Leverage linearization and Benders’ decomposition to convert mixed ILP problem into an ILP master problem and a linear programming subproblem, present a hybrid Quantum Benders’ decomposition algorithm Minimize total learning costs of CPU iterations QA Centralized
@@
TABLE XI: Details of QC Algorithms in Task Offloading
@@ IV-C2 Applications of Conventional QC for Task Offloading
Traditional task offloading, which involves scheduling function virtualization services to users, was explored in [182, 183]. Specifically, in [182], the network provided $K$ types of network function virtualization services, with virtual machines organized into groups dedicated to serving specific network functions. The scheduling problem for network function virtualization was initially framed as an Integer Linear Programming (ILP) challenge, subsequently converted into a QUBO model, and addressed by QA. Similarly, [183] addressed the task of offloading virtual network functions within a simple network setup. It provided a detailed implementation introduction, discussing the process of embedding the ILP problem into a QUBO format. It included the formation of the Q-matrix, crucial for solvability by QA, and the mapping of quantum hardware to the problem domain.
@@
With the evolution of 5G and beyond, edge computing emerges as a compelling paradigm aimed at delivering services to users by offloading computation tasks from users and cloud infrastructure to edge nodes situated in close proximity to users. In response to the challenge posed by limited edge server placement in IoV networks, a method called quantum edge server placement is developed in [185]. The method employed dual encoding, utilizing both binary and quantum encoding techniques, to represent the locations of edge servers. Leveraging the non-dominated sorting genetic algorithm, it effectively explored potential placement solutions. By employing an assessment function, it derived optimal placement schemes, thereby improving the efficiency of edge server deployment in IoV scenarios. However, it is worth noting that it did not account for network dynamics, such as the impact of vehicle mobility. QGA-based task offloading also found applications in intelligent agriculture, including irrigation and fertilization tasks [186]. Quantum operators utilized rotation gates and non-gates to update the matrix, mitigating the algorithm complexity by eliminating query operations from the rotation angle table during quantum rotation gate updates. However, it is worth noting that in practice, the latency requirements for drip irrigation and fertilization tasks were not stringent, and the demand for computational resources in these applications tended to be low. Consequently, these factors resulted in less attention being given to finding optimized solutions for such scenarios.
@@
Another conventional technique for task offloading is QPSO, widely applied in service placement within IoT networks [187], as well as in multi-task applications [189, 190, 191]. A QPSO method is proposed specifically for IoT task offloading in edge computing systems in [187]. The problem was formulated mathematically, with quantum particles efficiently designed to handle the entire IoT task offloading process. These particles guaranteed the generation of complete and valid solutions to the problem. Notably, the decoding process of quantum particles employed a novel double-hashing technique, ensuring uniform task offloading to the edge nodes. Furthermore, the fitness function was designed considering four distinct objectives: maximizing throughput, minimizing delay, reducing energy consumption, and balancing system load.
@@
Federated learning, a method of distributing computational tasks across multiple decentralized servers [184], can be considered a variant of task offloading. The approach addressed the challenge of training machine learning models across a distributed network without centralized data aggregation. The process involved formulating a joint participant selection and learning scheduling problem as a mixed ILP problem, aimed at minimizing the total learning costs. To tackle this optimization problem, a hybrid quantum-classical Benders’ Decomposition algorithm was proposed. By combining classical and QC methodologies, the approach aimed to efficiently solve the optimization problem associated with multi-modal federated learning in distributed networks.
@@
The previous studies fail to account for the impact of multiple tasks on system performance. For instance, in [187], while mathematical models involved multiple users with multiple tasks, there was no mutual influence among them. The task offloading problem was mathematically formulated as a mixed-integer nonlinear programming (MINLP) model in [188]. The formulation took into account constraints such as deadlines and resource capacities, with the objective of maximizing user satisfaction. Within the proposed QPSO algorithm, task scheduling solutions were mapped to positions in the search space using real number coding. Here, the value in each position denoted the computing resource to which the corresponding task is assigned. The algorithm adopted a population evolutionary approach of QPSO to iteratively search for the global best solution.
@@
The study presented in [189] introduced an innovative solution to the task offloading within MEC environments characterized by users and servers. In this study, the task offloading strategy was reformulated as an energy optimization problem. To address this challenge, an enhanced QPSO algorithm was proposed, aimed at efficiently solving the optimization problem for optimal task offloading. The effectiveness of the proposed QPSO-driven offloading strategy was demonstrated through its convergence. A bound state described by the probability density function, appeared at any interval of the completely searchable space with a certain probability. However, it was noteworthy that the impact of different tasks remained relatively weak, as multiple tasks from various users competed for the limited resources available at edge nodes. Additionally, the study lacked analysis regarding the impact of mobility within MEC environments. Similar limitations were observed in the study conducted by [190] within the IoV environment. Additionally, the assumption that the next hop was predetermined in scenarios where tasks were offloaded from vehicles to other vehicles lacks practicality. The first limitation was addressed in [191]. It identified two interrelated impacts among tasks: firstly, when a complex task arrived, it was offloaded to a coalition of edge nodes for collaborative resolution; secondly, the complex task was divided into multiple sub-tasks, each offloaded to distinct coalitions. The proposal calculated the Shapley value of the optimal solution by employing a similarity metric based on the task rewards of each combined edge group.

### v3

@@ III-D3 Applications of AI-driven QC for User Association
Several other studies [133, 134] integrate AI techniques with QC algorithms within their solutions while considering network dynamics. For example, [133] discusses the path planning for cellular-connected UAVs to minimize the weighted sum of time cost and expected outage duration. Traditional offline optimization techniques face challenges due to practical considerations like local building distribution and directional antenna radiation patterns. To overcome this, the authors propose a QDRL solution that maps the problem into MDP, enabling the UAV to find the optimal flying direction within each time slot. Unlike approaches focusing on designing multi-layer quantum networks or updating network parameters, this study proposes a quantum-inspired representation of experience prioritization to improve data efficiency. In practical applications, certain buffer experiences are sampled for training with undesirably high frequency, leading to over-fitting. Additionally, the limited buffer size exacerbates this issue, resulting in biased and unfair sampling performance. To address these challenges, experience prioritization is represented in terms of qubits and determined by linear coefficient quantum ground states. This quantum representation creates a connection between quantum eigenstates and the selection or rejection of specific experiences. By leveraging quantum amplitude amplification, the method enables the manipulation of quantum collapse to better manage experience sampling and prevent over-fitting. However, the analysis does not establish direct correlations among multiple UAVs, simplifying the model into individual path planning for each UAV.
Due to robust capacity of QDRL for addressing challenges including those characterized by a large state and action spaces, it can also be effectively applied in the optimization of multiple objectives. [134] addresses multiple interconnected challenges, encompassing UAV trajectory planning, NOMA user grouping, and power allocation. Within this framework, users within a group maintain communication with the UAV within their communication ranges along the UAV trajectory. The decision regarding energy consumption by the UAV directly influences the objective. Given the extensive and continuous nature of both action and state spaces, the paper introduces a QDRL approach to effectively tackling the complexity of the problem.
In conclusion, QC algorithms present a compelling approach for user association in wireless networks. Table VIII offers a consolidated overview of these studies. However, there are two open issues in these QC algorithms addressing user association. First, these works typically relied on computer simulations for QC algorithm implementation. Second, few studies account for the impact of mobility on performance, even in UAV networks.
@@ IV-A Overview
This section explores how QC principles can revolutionize various aspects of networking, including routing, task offloading, and content caching as illustrated in Fig. 14. These topics are crucial for optimizing network performance in scenarios with high mobility, dynamic environments, and complex resource management that are becoming increasingly important in the 6G era and beyond. Routing is a fundamental issue in wireless networking, as it forms the backbone for ensuring efficient data transmission. Task offloading and content caching, in turn, are often built upon effective routing strategies. By focusing on these areas, the paper addresses key research challenges that are vital for improving network efficiency, scalability, and reliability. Additionally, this focus emphasizes the potential of QC to tackle networking issues that are difficult to address with classical methods.
@@ IV-B1 Overview
Routing presents fundamental challenges in network systems, encompassing the task of determining optimal paths for data packets to travel from a source to a destination [136]. In wireless networks, where devices may be mobile and network conditions dynamically change, routing becomes even more complex. The primary goal of routing is to ensure efficient and reliable data transmission while minimizing latency, packet loss, and resource consumption. Existing approaches to routing encompass a variety of algorithms and protocols designed to address different aspects of routing optimization. Traditional methods like optimized link state routing protocol (OLSR) and ad hoc on-demand distance vector (AODV) attempt to dynamically establish and update routes based on network topology changes. However, they suffer from overhead due to frequent updates and incur latency in route discovery. Additionally, traditional routing algorithms may not adequately address challenges such as mobility, variable channel conditions, and energy constraints in wireless networks [137].
QC-driven approaches offer promising solutions to overcome the limitations of traditional routing methods. By leveraging the principles of quantum mechanics, they can explore vast solution spaces more efficiently and potentially find optimal or near-optimal routes in significantly less time. In addition, QC enables the development of AI-driven routing algorithms that can adapt and learn from network dynamics, optimizing routing decisions in real-time. These quantum-enhanced approaches have the potential to improve routing efficiency, minimize latency, and enhance network reliability.
TABLE IX: Summary of QC Works on Routing in Wireless Networks
@@
TABLE X: Details of QC Algorithms in Routing
@@ IV-B2 Applications of Conventional QC for Routing
The paper [139] introduces a multi-phase method to address vehicle routing with backhauls, mapping key elements of each phase onto QUBO models suitable for Ising machines. The two-phase method comprises a clustering phase followed by a routing phase, while the three-phase method includes a depot clustering phase, a vehicle clustering phase, and a routing phase. The clustering-related phases are expansions of the knapsack problem, while the routing phases are tackled as TSP. Its modeling process of QUBO formation can be regarded as a tutorial for applying QC in the field of networking.
@@
The paper [140] presents a multi-objective dynamic programming framework aimed at generating potentially Pareto-optimal routes by leveraging correlations among specific links, thereby significantly reducing the total number of routes. This framework extends the widely single-objective Bellman-Ford algorithm to handle multiple objectives. Additionally, the paper introduces a quantum-assisted algorithm that capitalizes on both the dynamic programming framework and the synergies between quantum programming and hybrid programming to effectively solve the multi-objective routing problem. Similar works from the authors with different objectives or different QC algorithms are presented in [156, 157]. The study [141], similar to [140], utilizes the Boyer-Brassard-Hoyer-Tapp quantum search algorithm, but in a different scenario within self-organizing networks. They introduce an optimal multi-objective quantum-assisted algorithm aimed at assessing legitimate routes while incorporating the concept of Pareto optimality to reduce complexity. However, the study overlooks the influence of node mobility, a critical characteristic in self-organizing networks like Ad-hoc networks.
@@
The energy optimization model introduced in [142] targets sensor-enabled IoT environments. A QC-driven solution is proposed for optimizing the problem, emphasizing quantum representation, measurement, and rotation angle. However, a notable drawback of this approach is that minimizing the total energy cost across all relaying paths does not necessarily equate to prolonging the network lifetime. There are other works [143] based on quantum ant colony algorithms to solve routing problems. It utilizes the metric, pheromone value inspired by ant colony algorithm, to measure a routing path. Due to dynamic network changes, it requires a large computation time to track pheromone traits, especially in a large wireless sensor network.
@@
Additional popular quantum evolution based proposals for routing issues include QGA solutions [144, 145] and QPSO solutions [146, 147, 148]. [144] improves the network robustness of connectivity for IoT topology duo to node failures caused by malicious attacks. It proposes a QGA algorithm to maximize the cumulative sum of the number of nodes in the maximum connected sub-graph of the network. Additionally, [145] is designed to be resilient, which adapts to any changes in a planned topology of deployed sinks in anticipation of sensor failures. Also it achieves energy consumption balance among sensors by equalizing the path length and brings down the cost of wireless sensor networks by minimizing the number of sinks. Multicast routing in wireless mesh networks are considered with the objective of minimizing network delay [146]. The routing problem is converted into an integer programming problem with QoS constraints and is solved by QPSO combined with loop deletion operation. Connectivity in routing is explored in [147, 148]. [147] considers heterogeneous scenarios of network topology for the optimal path configuration by exploring and exploiting the hunts. It applies multiple inputs from heterogeneous industrial IoT into QPSO techniques. The differential evolution operator and crossover operations are used for information interchange among the nodes to avoid trapping into local minima. Specifically, it studies the impact of p-degrees of network connectivity on network performance. The same scenario of IoT is also studied in [148] to optimize selection of cluster heads for multi-hop inter-cluster communication, which is different from user association in a single-cell communication in subsection III-D. It proposes a cooperative multiple-input-multiple-output connectivity scheme between sensors and cluster heads to extend network lifetime. QPSO is applied to select the optimum cooperative coalitions of each hop in the routing path.
@@
The preceding studies overlook node mobility, a critical aspect in multi-hop wireless networks where devices and nodes dynamically move in a network environment. Several works have employed QC algorithms to address the challenge. For instance, the optimized link state routing protocol, designed for mobile ad-hoc networks, is explored in [149, 150] from the same researchers, both leveraging QGA approaches. The dynamic topology inherent in mobile ad-hoc networks pose significant challenges in finding and sustaining an optimal end-to-end path. However, the heuristic Q-learning strategy in RL enables dynamic adjustment of routing paths through interactions with environments. An enhanced QEA strategy has incorporated a novel Q-learning strategy, optimizing the selection of multi-point relays [149]. This improvement addresses the limitations of traditional protocols and demonstrates properties of convergence and global optimization. Other network aspects are additionally considered in [150], such as the delivery rate of data packets, and overhead of network topology control.
Additional networks, including UAV networks, are examined with a focus on node mobility in [151, 152], utilizing QPSO and quantum ant lion algorithm, respectively. Specifically, [151] combines the differential evolution algorithm with QPSO to further enhance the performance of both algorithms. The primary objective is to describe a general route planning strategy. A route planner for UAVs is designed to generate a safe and flyable path in the presence of different threat environments. [152] presents an approach for designing an energy-efficient neuro-fuzzy cluster-based topology construction with a metaheuristic route planning algorithm. The model utilizes three input parameters: residual energy in UAVs, average distance to nearby UAVs, and UAV degree, for selecting cluster heads and constructing the topology. Additionally, quantum ant lion based metaheuristic route planning is employed to select an optimal set of routes for inter-cluster UAV communication. However, the proposal divides the problem into two independent sub-problems, potentially leading to sub-optimal solutions.
@@ IV-B3 Applications of AI-driven QC for Routing
In recent years, AI techniques, such as RL and NNs, can be integrated with QC to enhance decision-making. AI-driven QC has emerged as a promising paradigm for addressing routing challenges in complex wireless environments.
@@
A framework [153] is introduced aiming at leveraging QRL to enhance the design of next-generation wireless systems. The framework, termed QRL for relay and transmit power selection, enables sensor nodes to make optimal decisions regarding relay and transmit power selection by leveraging both present and past local state knowledge. Initially, the relay and transmit power selection problem is formulated as MDP. Subsequently, the QRL optimization aspect of the MDP is addressed, focusing on jointly optimizing energy consumption and throughput to maximize network utility. To solve this MDP problem, a QRL algorithm is developed, drawing upon Grover’s iteration. The algorithm enables efficient decision-making within the network, facilitating enhanced performance and resource utilization. Especially, the simulation is implemented using quantum virtual machine hosted on Rigetti’s forest platform, which provides an additional approach to validating QC algorithms. QMARL is also applied in tackling routing problems. The work [154] introduces a quantum-based solution for MARL tailored for autonomous control and coordination of multiple robots within smart factory environments. The delivery problem shares similarities with the routing problem. It presents an innovative QMARL framework, featuring centralized training and decentralized execution. The framework leverages parameter sharing on policy, design of variational quantum circuits, and 2-variables dense encoding. In particular, the proposed system enables the QDRL server to make distributed and sequential decisions for each autonomous mobile robot. These decisions revolve around determining the optimal allocation of goods, including the number of goods to carry in each autonomous mobile robot and when to request quality control. The objective is to prevent both overflow and underflow scenarios during the delivery of goods by each autonomous mobile robot.
@@
The previous studies focusing on AI-driven QC overlook the consideration of node mobility in wireless networks. This aspect is addressed by a limited number of studies, such as [155]. Multiple mobile UAV platforms act as mobile BSs to gather data from other devices. Note that trajectory planning of multiple UAVs falls within the domain of routing, as data from devices are relayed by multiple UAVs to reach the destination. The proposed QMARL-based algorithm presents a significant advantage in reducing action dimensionality. By utilizing projective value measures, it efficiently compresses the dimensionality to a logarithmic scale. This dimensionality reduction facilitates improved convergence during training, particularly in large-scale MARL scenarios.
@@ IV-C1 Overview
Research in task offloading encompasses a spectrum of crucial topics [158, 159]. Efficient task partitioning and scheduling, ensuring optimal allocation of tasks between edge devices and the cloud, form a foundational aspect. The effective management of limited edge resources from computing, storage, to communication, is a critical consideration [160].
TABLE XI: Summary of QC Works on Task Offloading in Wireless Networks
@@
[163] A computing graph network with vertices as computing servers, and edges as vertex connection links Leverage linearization and Benders’ decomposition to convert mixed ILP problem into an ILP master problem and a linear programming sub-problem, present a hybrid Quantum Benders’ decomposition algorithm Minimize total learning costs of CPU iterations QA Centralized
@@
TABLE XII: Details of QC Algorithms in Task Offloading
@@ IV-C2 Applications of Conventional QC for Task Offloading
Traditional task offloading, which involves scheduling function virtualization services to users, is explored in [161, 162]. Specifically, in [161], the network provides $K$ types of network function virtualization services, with virtual machines organized into groups dedicated to serving specific network functions. The scheduling problem for network function virtualization is initially framed as an integer linear programming (ILP) challenge, subsequently converted into a QUBO model, and addressed by QA. Similarly, [162] addresses the task of offloading virtual network functions within a simple network setup. It provides a detailed implementation introduction, discussing the process of embedding the ILP problem into a QUBO format. It includes the formation of the Q-matrix, crucial for solvability by QA, and the mapping of quantum hardware to the problem domain.
@@
With the evolution of 5G/6G and beyond, edge computing emerges as a compelling paradigm aimed at delivering services to users by offloading computation tasks from users and cloud infrastructure to edge nodes situated in close proximity to users. In response to the challenge posed by limited edge server placement in IoV networks, a method called quantum edge server placement is developed in [164]. The method employs dual encoding, utilizing both binary and quantum encoding techniques, to represent the locations of edge servers. Leveraging the non-dominated sorting genetic algorithm, it effectively explores potential placement solutions. By employing an assessment function, it derives optimal placement schemes, thereby improving the efficiency of edge server deployment in IoV scenarios. However, it is worth noting that it does not account for network dynamics, such as the impact of vehicle mobility. QGA-based task offloading also finds applications in intelligent agriculture, including irrigation and fertilization tasks [165]. Quantum operators utilizes rotation gates and non-gates to update the matrix, mitigating the algorithm complexity by eliminating query operations from the rotation angle table during quantum rotation gate updates. However, it is worth noting that in practice, the latency requirements for drip irrigation and fertilization tasks are not stringent, and the demand for computational resources in these applications tends to be low. Consequently, these factors result in less attention being given to finding optimized solutions for such scenarios.
@@
Another conventional technique for task offloading is QPSO, widely applied in service placement within IoT networks [166], as well as in multi-task applications [168, 169, 170]. A QPSO method is proposed specifically for IoT task offloading in edge computing systems in [166]. The problem is formulated mathematically, with quantum particles efficiently designed to handle the entire IoT task offloading process. These particles guarantee the generation of complete and valid solutions to the problem. Notably, the decoding process of quantum particles employs a novel double-hashing technique, ensuring uniform task offloading to the edge nodes. Furthermore, the fitness function is designed considering four distinct objectives: maximizing throughput, minimizing delay, reducing energy consumption, and balancing system load.
@@
Federated learning, a method of distributing computational tasks across multiple decentralized servers [163], can be considered a variant of task offloading. The approach addresses the challenge of training ML models across a distributed network without centralized data aggregation. The process involves formulating a joint participant selection and learning scheduling problem as a mixed ILP problem, aimed at minimizing the total learning costs. To tackle this optimization problem, a hybrid quantum-classical benders’ decomposition algorithm is proposed. By combining classical and QC methodologies, the approach aims to efficiently solve the optimization problem associated with multi-modal federated learning in distributed networks.
@@
The previous studies fail to account for the impact of multiple tasks on system performance. For instance, in [166], while mathematical models involve multiple users with multiple tasks, there is no mutual influence among them. The task offloading problem is mathematically formulated as a mixed-integer nonlinear programming model in [167]. The formulation takes into account constraints such as deadlines and resource capacities, with the objective of maximizing user satisfaction. Within the proposed QPSO algorithm, task scheduling solutions are mapped to positions in the search space using real number coding. Here, the value in each position denotes the computing resource to which the corresponding task is assigned. The algorithm adopts a population evolutionary approach of QPSO to iteratively search for the global best solution.
@@
The study presented in [168] introduces an innovative solution to task offloading within mobile edge computing (MEC) environments characterized by users and servers. In this study, the task offloading strategy is reformulated as an energy optimization problem. To address this challenge, an enhanced QPSO algorithm is proposed. Conventional PSO approaches constrain the search range of particles, restricting them from exploring positions that are far from the swarm, even if those positions yield better solutions than the current best. This study introduces a novel approach by utilizing quantum state representation to address this limitation. By representing multiple potential solutions as quantum states before measurement, the method enables particles to explore a broader solution space, thereby reducing the likelihood of getting trapped in local optima. The effectiveness of the proposed QPSO-driven offloading strategy is demonstrated through its convergence. A bound state described by the probability density function, appears at any interval of the completely searchable space with a certain probability. However, it is noteworthy that the impact of different tasks remains relatively weak, as multiple tasks from various users compete for the limited resources available at edge nodes. Additionally, the study lacks analysis regarding the impact of mobility within MEC environments. Similar limitations are observed in the study conducted by [169] within the IoV environment. Additionally, the assumption that the next hop is predetermined in scenarios where tasks are offloaded from vehicles to other vehicles lacks practicality. The first limitation is addressed in [170]. It identifies two interrelated impacts among tasks: firstly, when a complex task arrives, it is offloaded to a coalition of edge nodes for collaborative resolution; secondly, the complex task is divided into multiple sub-tasks, each offloaded to distinct coalitions. The proposal calculates the Shapley value of the optimal solution by employing a similarity metric based on task rewards of each combined edge group.